OpenStudy (fancysharkk):

What is each number in standard form? Drag the answer into the box to match each number. 8.38×10108.38×1010 ​ 8.308×1068.308×106 ​ ​ 8.038×1088.038×108 ​ 8,308,000,000803,800,000,00083,800,000,000803,800,0008,308,0008,380,000,000,000

OpenStudy (fancysharkk):

Uhh i have to take a screenshot..

OpenStudy (jiteshmeghwal9):

\[8.38 * 10^{10} = 838 * 10^{-2}*10^{10}\]can u solve it further ?

OpenStudy (fancysharkk):

Uhh. this is really hard on me and i cant get a bad grade on it.but i can try?

OpenStudy (frozen15):

So what do you think? Or do you not know as for what your saying you cant do this? ^

OpenStudy (fancysharkk):

Im trying to figure it out. i got 1 try on this test and i dont wanna fail :~:

OpenStudy (jiteshmeghwal9):

\[838 * 10^8\]so place 8 zeros on the back of 838

OpenStudy (jiteshmeghwal9):

@FancySharkk

OpenStudy (fancysharkk):

i think i did it

OpenStudy (frozen15):

I understand , I'm well focused on trying to get good grades and not fail more than once on a test myself haha , are you using k12?

OpenStudy (fancysharkk):

yeah. i am using k12

OpenStudy (fancysharkk):

its just last year i failed miserably so i dont want that to happen again and probably get kicked out of k12

OpenStudy (jiteshmeghwal9):

while solving them u did not notice the decimal places, did u ?

OpenStudy (fancysharkk):

i did notince the decimal places

OpenStudy (andrewishere):

What is up with this problem, like, for real. Everyone in here already explained it well but, for real.

OpenStudy (fancysharkk):

what?

OpenStudy (jiteshmeghwal9):

\(8.38*10^{10}\) =\(838*10^{-2}*10^{10}\) =\(838*10^8\) now u only need is to place 8 zeros on the back of 838

OpenStudy (jiteshmeghwal9):

83800000000

OpenStudy (frozen15):

@FancySharkk I am using the same program and I doubt they will kick a student. If your constantly not paying attention to the questions and facts , they will have you kicked or just restart the whole semester or grade. You got to be careful. I never had this problem because I never failed my classes , I am pretty high ranked on my reading level. Not bragging or anything but try using solutions the lessons/units give you , that's why they are their , to help you. If you have any hard troubles , if you have a teacher , which I don't , but if you do try having your guardian contact your teacher for more help such as practice sheets , more equations , etc. Just stay focused :)

OpenStudy (fancysharkk):

ok ^u^

OpenStudy (frozen15):

:D if you need any further help , let me know but for right now I cant cause I am busy doing my essay on word so good luck ^.^